# Neural density functional theory of liquid-gas phase coexistence

### v1

@@ III.6 Inhomogeneous fluids: predictions for drying, capillary evaporation and local fluctuations
Figure 9: Density $\rho(x)$ and local compressibility $\chi_{\mu}(x)$ profiles for drying at a hard wall (a, b) and for capillary evaporation between hard walls with weak attraction (c, d), cf. Eq. (28). Results are shown for chemical potentials $\mu=\mu_{\mathrm{coex}}+\Delta\mu$ approaching, from the liquid side, the respective bulk coexistence values $\mu_{\mathrm{coex}}$—see text. In the drying case $k_{B}T/\varepsilon=0.93$ and density profiles (a) show a gas layer at the wall whose thickness grows for decreasing values of $\Delta\mu$. The local compressibility profile (b) develops a sharp maximum (note the logarithmic scale) at a position located in the emerging gas-liquid interface whose height grows rapidly on approaching bulk coexistence. In the case of capillary evaporation $k_{B}T/\varepsilon=1.0$ and the density profiles in (c) refer to the condensed ‘liquid’ states (solid lines) and the stable ’gas’ states (dotted lines) plotted for the same chemical potential differences as in (a). The inset displays the difference $\Delta\Omega=\Omega_{l}-\Omega_{g}$ between the liquid and gas grand potentials $\Omega_{l}$ and $\Omega_{g}$ per lateral system area $A$. Positive values of $\Delta\Omega$ indicate metastability of the condensed state. Similar to (b), the local compressibility profiles (d) display maxima located in the interface which increase for decreasing $\Delta\mu$.
Previous subsections have focused mainly on bulk behaviour, i.e. the fluid in the absence of any external potential. Here we turn to inhomogeneous systems and, in particular, to the physics of adsorption at substrates and of fluids in confinement. Specifically, we show that the neural functional allows one to capture accurately fluid structure for phenomena associated with phase transitions at state points close to bulk coexistence. These are challenging to access via simulation studies. We investigate the phenomena of i) density depletion and drying at a hard-wall and ii) capillary evaporation in a planar slit. Quantifying the properties associated with such phenomena serves as a stringent test of the neural functional; recall that the training data consist of systems with randomized state points and inhomogeneities not tailored to feature the subtle physics occurring near surface phase transitions.
@@
The parameter $\varepsilon_{w}$ sets the attraction strength and the minimum of the potential well at $x_{m}=(2/5)^{1/6}\sigma$ is shifted to $x=0$. The right wall is obtained analogously by a mirrored version of Eq. (28) and the symmetric confinement potential is the sum of these. We choose this potential to correspond to a recent MC simulation study by Wilding et al. [72] and set the attraction strength of the wall $\varepsilon_{w}=0.01\varepsilon$, wall separation $L=25\sigma$ and temperature $k_{B}T/\varepsilon=1.0$ to be the same as in Ref. [72]. For such a weak attraction we expect complete drying to occur at a single wall, i.e. the corresponding contact angle should be $180^{\circ}$. Results for the density profiles are shown in Fig. 9(c) for the same values of $\Delta\mu$ as in panel (a). For this temperature the neural prediction is $\mu_{\mathrm{coex}}/\varepsilon\approx-3.46154$. The density profiles are typical of those found for a fluid in a slit with hard or very weakly attractive walls: a liquid-like density plateau forms at the center of the slit and as $\Delta\mu$ decreases increasing density depletion occurs and eventually a layer of gas develops at the walls. An early (analytical) DFT study [73], see also Ref. [74], showed that the ‘liquid’ becomes metastable with respect to the evaporated ‘gas’, whose density is small throughout the slit, and determined the capillary evaporation (first-order) transition point for a model fluid.
@@
The contact values $\chi_{\mu}(0^{+})$ at the hard wall remain almost unchanged as $\Delta\mu$ is reduced. This behaviour is consistent with the contact theorem $\chi_{\mu}(0^{+})=\beta\rho_{b}$ that establishes an important connection between the bulk density $\rho_{b}$ far from the wall and a quantity local to the wall [75, 77, 78]. Our numerical results satisfy this sum rule to within $1\%$, apart from the lowest value of $\Delta\mu$, where numerical errors become more important. In our investigation of capillary evaporation we chose our system to correspond to that of the MC simulations in Ref. [72] which examined the local compressibility at bulk coexistence. The density and local compressibility profiles that we obtain mimic those observed in simulation. However, there are subtleties in making direct comparisons that we return to in Sec. IV.2. We chose to study the profiles at decreasing values of $\Delta\mu$ monitoring the growth of the maximum in the local compressibility which increases in a similar fashion to what is observed for the complete drying case at a single hard wall in Fig. 9(b).
@@ IV.1 Methodology
Employing the neural functional approach is very different from what one encounters in pure simulation studies or in analytical density functional studies where an explicit approximation to the excess free energy functional is provided from the outset. Setting up a neural functional from scratch requires an upfront investment in compute resources in order to deliver a suitable training data set. This (perceived) hurdle is not encountered in working with simulations only; the first run can deliver useful information. However, once the initial threshold has been passed the method scales very well. As re-training the neural network itself is relatively cheap, it is entirely feasible and sensible to provide additional training data, be it for changing system size $1/L$, extending parameter ranges, such as temperature $T$, or simply to provide better statistics in an iterative cycle, possibly guided by active machine learning techniques. Crucially, the initial investment is never lost, as the original training data can continue to be used in updated training cycles.
@@
What sets aside our neural functional approach is the range of phenomena and results that it can describe: these far exceed the information provided during training. How this comes about is non-trivial and relies upon the mathematical structure of DFT. Investigating physical phenomena, i.e. calculating the structure and thermodynamic properties, reduces to standard analysis tasks within the (neural-network-based) functional mapping $c_{1}(\mathbf{r};[\rho],T)$. Inputting only MC training data of one-body profiles in planar geometry and then examining $c_{1}(\mathbf{r};[\rho],T)$ through the functional lens provides access to quantities which could not be obtained directly from the input data. Indeed determining these usually requires advanced simulation techniques [9]; note the array of computational methods tailored to study liquid-gas coexistence, interfacial profiles, free energies and the surface tension as well as two-body correlation functions. These properties and the physical scenarios they encompass follow, in our treatment, from a single numerical object, namely the neural network representing $c_{1}(\mathbf{r};[\rho],T)$. Our results achieve accuracy comparable with direct simulations of phase coexistence but with much reduced computational cost, and without requiring any additional simulations after training.
@@
We consider our present methodology to be more closely interwoven with the basic theoretical physics at play, in this case DFT, than are the more generic data-based machine learning methods [1, 2, 3, 4]. Importantly, the formal structure of density functional relationships between generating functionals and hierarchial levels of correlation functions is built into our neural density functional approach and guides application using functional calculus (see Sec. II.4). In this regard, it might be interesting to explore similarities to and differences from other recent work [80, 81, 82] which employs liquid integral equation theory in the context of physics-informed machine learning.
@@ IV.2 Physical phenomena investigated
In this subsection we summarize some of the physics that we investigated together with the results we found and raise issues that require further investigation. For the bulk LJ (continuum) fluid the trained neural direct correlation functional yields an accurate description of pair structure. The pair correlation function $g(r)$ and static structure factor $S(k)$ exhibit signatures of fluids with attractive interactions, see Fig. 2 (Sec. III.1). Moreover, a comprehensive account of very subtle correlation phenomena can be obtained with relative computational ease from our determination of the bulk pair direct correlation function. Specifically we determined the lines in the phase diagram where the isothermal compressibility and true correlation length have their maximal value, and the Fisher-Widom line that denotes the crossover from damped oscillatory to purely monotonic decay of $g(r)$ at large distances, see Fig. 3 (Sec. III.2). Recall, once again, that the neural network had not encountered bulk pair correlation functions, let alone their asymptotic decay, in training.
@@
Our results for the spinodal and the location of the liquid-gas critical point as determined from the divergence of the true correlation length, or equivalently of the compressiblity, are also obtained from $c_{2}^{b}(r)$; these are shown in Fig. 3. The OZ plots in Fig. 8 provide further information about the approach to criticality from above $T_{c}$: we find mean-field critical exponents for the OZ correlation length and for the compressibility. Separately we find that the true correlation length diverges in the same fashion as the OZ correlation length. The spinodal is, of course, a mean-field concept; in reality there is only a horizontal tie-line in the density-temperature plane linking coexisting gas and liquid. Analytical DFT approximations that include attraction predict a spinodal. It seems that our neural functional generates $c_{2}^{b}(r)$, for appropriate densities and temperatures, with the form that gives rise to a spinodal. We examine this further in the next subsection.
@@
In many respects our determination of the gas-liquid binodal (bulk coexistence curve) is one of the most striking results; see Sec. III.3. Our training data had no knowledge of bulk phase separation. Comparison with existing simulation results in Fig. 4 shows how well the neural prediction performs and how it out-performs the standard analytical DFT. Using fitting procedures applied previously to simulation data [58, 57] for coexisting densities we found evidence for a non-classical critical exponent $\beta$ and obtained an estimate of $k_{B}T_{c}/\varepsilon$ that is about one percent lower than we obtained from our analyses approaching $T_{c}$ from above. The reasons for the differences between sub- and super-near-critical behaviour remain to be ascertained. These are likely to be associated with subtle finite size effects. We have demonstrated that the neural prediction for the binodal is i) highly accurate and ii) internally consistent: results from neural density functional minimization of coexistence states that feature both gas and liquid in a single computational system agree with those from functional line integration for the free energy. Practical implementation of the latter method requires only a cheap numerical routine; the neural functional is the integrand in Eq. (10). The bulk equation of state from functional line integration displays a van der Waals loop, see Fig. 6, at subcritical temperatures. We comment further on this result in Sec. IV.3.
@@
The former method of direct numerical stabilization of phase coexistence provides access to one-body liquid-gas interfacial structure. As remarked in Sec. III.4, the density profiles $\rho(x)$ for the two lowest temperatures show the presence of damped oscillations extending into the liquid, see Fig. 7(a). Such behaviour is commonly found in DFT calculations that utilize the standard mean-field approximation (4), e.g. Ref. [67] for the square-well and Tschopp et al. [83] for the hard-core Yukawa model. Equivalent behaviour is also found in FMT-based DFT calculations for the Asakura-Oosawa model that describes colloid-polymer mixtures. Pronounced oscillations are found on the (colloid-rich) liquid side of the fluid-fluid interface for states in the neighbourhood of the triple point [84]. The physical origin of the oscillations lies, of course, in the packing of particles in the dense liquid. The argument [67], based on asymptotics, is that in the oscillatory region of the bulk phase diagram, as determined by the Fisher-Widom line, the one-body liquid-gas density profile should decay into the bulk liquid at coexistence with the same exponential decay length $1/\alpha_{0}$ and same wavelength $2\pi/\alpha_{1}$ as would be determined for the bulk state from the pole analysis, i.e. Eqs. (17)–(20). Unfortunately there is no simple means of determining the amplitude of oscillations. Moreover, the argument is based on mean-field ideas and omits effects of thermally induced capillary wave fluctuations. The physical picture adopted in Ref. [67], and in subsequent work, is that the DFT results provide a ‘bare’ profile that is then ‘dressed’ by unfreezing the fluctuations which serve to erode the oscillations. For a Gaussian treatment of the fluctuations the decay length and the wavelength of the oscillations are unchanged but the amplitude is reduced by a factor that depends on the interfacial roughness which depends, in turn, upon the interfacial area $A$; see Brader et al. [85] for a thorough discussion. The profiles we find here for the LJ fluid exhibit very weak oscillations even at the lowest temperature shown, suggesting that our neural functional might already capture some fluctuation effects. In this context, it is instructive to take note of Tschopp et al. [83] who show that profiles obtained from a DFT that incorporates correlations into a reference hard-sphere system yields liquid-gas density profiles that have much less pronounced oscillations than those that emerge from employing Eq. (4). It is tempting to surmise that our neural functional provides a very accurate, albeit still mean-field, description of the ‘bare’ interface.
@@
The liquid-gas surface tension is not beset by subtle issues of including capillary wave fluctuations and we expect our results, obtained from an inhomogeneous functional line integral, to be reliable. Indeed these match closely high quality simulation data for the LJ system, see Fig. 7(b).
@@ IV.3 Discovering phase coexistence and extrapolation of functional mappings
The successful reproduction of results with supercritical training only, cf. Sec. III.7, calls for a reassessment of the importance of the one-body direct correlation functional $c_{1}(\mathbf{r};[\rho],T)$; recall this is the quantity we seek to capture from simulation data via a neural network. It seems feasible to infer this particular functional mapping even when excluding substantial ranges of the parameter space in the training data. Our findings point to the fundamental nature of the one-body direct correlation functional and its favorable mathematical properties. It is arguably the object most appropriate for machine learning tasks—recall also the simplicity of Eq. (7) for determining $c_{1}(x)$ from simulation data—in contrast to other possible functional mappings [21, 22]. Importantly our results show that liquid-gas phase coexistence is an emerging phenomenon that can already be gleaned far from its onset, i.e. from training data taken above the critical temperature.
@@
We note in this context that the extrapolation capabilities of the neural functional, and of the underlying functional mapping it represents, are arguably much more extensive than one might initially expect. Recall that the data used during training, see Sec. II.3, comprise only true equilibrium profiles obtained via many-body simulations, and that the resulting density profiles fulfill the Euler-Lagrange equation (8) by construction. However, when using the trained neural functional for predictions, one cannot guarantee a priori that only such ‘physical’ density profiles are encountered. This may be the case during the iteration of Eq. (8) as well as for the evaluation of the excess free energy via the parametrized functional integral (10). For the latter the scaled density profiles $\rho_{a}(\mathbf{r})=a\rho(\mathbf{r})$ are not merely transient, rather they enter as genuine contributions to the total value of $F_{\mathrm{exc}}([\rho],T)$.
@@
That the predictions remain accurate despite requiring the evaluation of the neural functional with possibly ‘unphysical’ density profiles, i.e. ones that cannot correspond to minimization in the presence of any external potential, indicates there is a well-defined continuation of the functional relationship to such states. Furthermore, it seems feasible to infer this extended mapping from physical data alone, using only the restricted function space of the true equilibrium density profiles that provided input. We attribute the successful extrapolation to i) the beneficial mathematical properties of the functional mapping $c_{1}(\mathbf{r};[\rho],T)$ and ii) the prowess of the utilized machine learning techniques.
In this regard, we note that the van der Waals loop in Fig. 6 takes on a different status from that usually assumed: it arises here from the neural prediction for metastable and unstable bulk densities, invoking explicitly the functional extrapolation to such states. This is also pertinent for the determination of the surface tension, cf. Fig. 7, where one includes intermediate density profiles in the functional line integration of the liquid-gas interface in evaluating $\gamma_{lg}$. Similarly, bulk pair direct correlation functions are obtained within the coexistence region and even in the interior of the spinodal from the evaluation of $c_{1}(\mathbf{r};[\rho],T)$ and its functional derivative (9). This allows us to determine related quantities such as the isothermal compressibility $\chi_{T}$ and the correlation length $\xi$, see Fig. 3, in states of the bulk fluid that might not be physically accessible. Of course, this is not different from using analytical DFT: an explicit approximation for the excess free energy functional will provide $c_{2}^{b}(r)$ for all states.

### v2

@@ III.6 Inhomogeneous fluids: predictions for drying, capillary evaporation and local fluctuations
Figure 9: Density $\rho(x)$ and local compressibility $\chi_{\mu}(x)$ profiles for drying at a hard wall (a, b) and for capillary evaporation between hard walls with weak attraction (c, d), cf. Eq. (28). Results are shown for chemical potentials $\mu=\mu_{\mathrm{coex}}+\Delta\mu$ approaching, from the liquid side, the respective bulk coexistence values $\mu_{\mathrm{coex}}$—see text. We investigate $\Delta\mu/\varepsilon=0.3,0.1,0.01,0.001$, and $0.0001$, as labeled. In the drying case $k_{B}T/\varepsilon=0.93$ and density profiles (a) show a gas layer at the wall whose thickness grows for decreasing values of $\Delta\mu$. The local compressibility profile (b) develops a sharp maximum (note the logarithmic scale) at a position located in the emerging gas-liquid interface whose height grows rapidly on approaching bulk coexistence. In the case of capillary evaporation $k_{B}T/\varepsilon=1.0$ and the density profiles in (c) refer to the condensed ‘liquid’ states (solid lines) and the stable ‘gas’ states (dotted lines) plotted for the same chemical potential differences as in (a). The inset displays the difference $\Delta\Omega=\Omega_{l}-\Omega_{g}$ between the liquid and gas grand potentials $\Omega_{l}$ and $\Omega_{g}$ per lateral system area $A$. Positive values of $\Delta\Omega$ indicate metastability of the condensed state. Similar to (b), the local compressibility profiles (d) display maxima located in the interface which increase for decreasing $\Delta\mu$.
Previous subsections have focused mainly on bulk behavior, i.e. the fluid in the absence of any external potential. Here we turn to inhomogeneous systems and, in particular, to the physics of adsorption at substrates and of fluids in confinement. Specifically, we show that the neural functional allows one to capture accurately fluid structure for phenomena associated with phase transitions at state points close to bulk coexistence. These are challenging to access via simulation studies. We investigate the phenomena of i) density depletion and drying at a hard-wall and ii) capillary evaporation in a planar slit. Quantifying the properties associated with such phenomena serves as a stringent test of the neural functional; recall that the training data consist of systems with randomized state points and inhomogeneities not tailored to feature the subtle physics occurring near surface phase transitions.
@@
The parameter $\varepsilon_{w}$ sets the attraction strength and the minimum of the potential well is shifted to $x=0$ by setting $x_{m}=(2/5)^{1/6}\sigma$. The right wall is obtained analogously by a mirrored version of Eq. (28) and the symmetric confinement potential is the sum of these. We choose this potential to correspond to a recent MC simulation study by Wilding et al. [74] and set the attraction strength of the wall $\varepsilon_{w}=0.01\varepsilon$, wall separation $L=25\sigma$ and temperature $k_{B}T/\varepsilon=1.0$ to be the same as in Ref. [74]. For such a weak attraction we expect complete drying to occur at a single wall, i.e. the corresponding contact angle should be $180^{\circ}$. Results for the density profiles are shown in Fig. 9(c) for the same values of $\Delta\mu$ as in panel (a). For this temperature the neural prediction is $\mu_{\mathrm{coex}}/\varepsilon\approx-3.46154$. The density profiles are typical of those found for a fluid in a slit with hard or very weakly attractive walls: a liquid-like density plateau forms at the center of the slit and as $\Delta\mu$ decreases increasing density depletion occurs and eventually a layer of gas develops at the walls. An early (analytical) DFT study [75], see also Ref. [76], showed that the ‘liquid’ becomes metastable with respect to the evaporated ‘gas’, whose density is small throughout the slit, and determined the capillary evaporation (first-order) transition point for a model fluid.
@@
The contact values $\chi_{\mu}(0^{+})$ at the hard wall remain almost unchanged as $\Delta\mu$ is reduced. This behavior is consistent with the contact theorem $\chi_{\mu}(0^{+})=\beta\rho_{b}$ that establishes an important connection between the bulk density $\rho_{b}$ far from the wall and a quantity local to the wall [77, 79, 80]. Our numerical results satisfy this sum rule to within $1\%$, apart from the lowest value of $\Delta\mu$, where numerical errors become more important. In our investigation of capillary evaporation we chose our system to correspond to that of the MC simulations in Ref. [74] which examined the local compressibility at bulk coexistence. The density and local compressibility profiles that we obtain mimic those observed in simulation. However, there are subtleties in making direct comparisons that we return to in Sec. IV.2. We chose to study the profiles at decreasing values of $\Delta\mu$ monitoring the growth of the maximum in the local compressibility which increases in a similar fashion to what is observed for the complete drying case at a single hard wall in Fig. 9(b).
@@ IV.1 Methodology
Employing the neural functional approach is very different from what one encounters in pure simulation studies or in analytical density functional studies where an explicit approximation to the excess free energy functional is provided from the outset. Setting up a neural functional from scratch requires an upfront investment in compute resources in order to deliver a suitable training data set. This (perceived) hurdle is not encountered in working with simulations only; the first run can deliver useful information. However, once the initial threshold has been passed the method scales very well. As retraining the neural network itself is relatively cheap, it is entirely feasible and sensible to provide additional training data, be it for changing system size $1/L$, extending parameter ranges, such as temperature $T$, or simply to provide better statistics in an iterative cycle, possibly guided by active machine-learning techniques. Crucially, the initial investment is never lost, as the original training data can continue to be used in updated training cycles.
@@
What sets aside our neural functional approach is the range of phenomena and results that it can describe: these far exceed the information provided during training. How this comes about is nontrivial and relies upon the mathematical structure of DFT. Investigating physical phenomena, i.e. calculating the structure and thermodynamic properties, reduces to standard analysis tasks within the (neural-network-based) functional mapping $c_{1}(\mathbf{r};[\rho],T)$. Inputting only MC training data of one-body profiles in planar geometry and then examining $c_{1}(\mathbf{r};[\rho],T)$ through the functional lens provides access to quantities which could not be obtained directly from the input data. Indeed determining these usually requires advanced simulation techniques [9]; note the array of computational methods tailored to study liquid-gas coexistence, interfacial profiles, free energies and the surface tension as well as two-body correlation functions. These properties and the physical scenarios they encompass follow, in our treatment, from a single numerical object, namely the neural network representing $c_{1}(\mathbf{r};[\rho],T)$. Our results achieve accuracy comparable with direct simulations of phase coexistence but with much reduced computational cost, and without requiring any additional simulations after training.
@@
We consider our present methodology to be more closely interwoven with the basic theoretical physics at play, in this case DFT, than are the more generic data-based machine-learning methods [1, 2, 3, 4]. Importantly, the formal structure of density functional relationships between generating functionals and hierarchical levels of correlation functions is built into our neural density functional approach and guides application using functional calculus (see Sec. II.4). In this regard, it might be interesting to explore similarities to and differences from other recent work [82, 83, 84] which employs liquid integral equation theory in the context of physics-informed machine learning.
@@ IV.2 Physical phenomena investigated
In this subsection we summarize some of the physics that we considered together with the results we found and raise issues that require further investigation. For the bulk LJ (continuum) fluid the trained neural direct correlation functional yields an accurate description of pair structure. The pair correlation function $g(r)$ and static structure factor $S(k)$ exhibit signatures typical for fluids with attractive interactions, see Fig. 2 (Sec. III.1). Moreover, a comprehensive account of very subtle correlation phenomena can be obtained with relative computational ease from our determination of the bulk pair direct correlation function. Specifically we determined the lines in the phase diagram where the isothermal compressibility and true correlation length have their maximal value, and the Fisher-Widom line that denotes the crossover from damped oscillatory to purely monotonic decay of $g(r)$ at large distances, see Fig. 3 (Sec. III.2). Recall, once again, that the neural network had not encountered bulk pair correlation functions, let alone their asymptotic decay, in training.
@@
Our results for the spinodal and the location of the liquid-gas critical point as determined from the divergence of the true correlation length, or equivalently of the compressiblity, are also obtained from $c_{2}^{b}(r)$; these are shown in Fig. 3. The OZ plots in Fig. 8 provide further information about the approach to criticality from above $T_{c}$: we find mean-field critical exponents for the OZ correlation length and for the compressibility. Separately we find that the true correlation length diverges in the same fashion as the OZ correlation length. The spinodal is a concept commonly associated with employing analytic functionals that incorporate attraction, e.g. via the mean-field approximation (4); of course, in reality there is only a horizontal tie-line in the density-temperature plane linking coexisting gas and liquid. It seems that our simulation-based machine-learning procedure generates a neural functional, which, for appropriate bulk densities and temperatures, also gives rise to a spinodal. We examine this further in the next subsection.
@@
In many respects our determination of the gas-liquid binodal (bulk coexistence curve) is one of the most striking results; see Sec. III.3. Our training data had no knowledge of bulk phase separation. Comparison with existing simulation results in Fig. 4 shows how well the neural prediction performs and how it out-performs the standard analytical DFT. Using fitting procedures applied previously to simulation data [60, 59] for coexisting densities we found evidence for a nonclassical critical exponent $\beta$ and obtained an estimate of $T_{c}$ that agrees very closely with the best simulation estimate. The subtleties involved with finite-size effects were discussed in Sec. III.5. We have demonstrated that the neural prediction for the binodal is i) highly accurate and ii) internally consistent: results from neural density functional minimization of coexistence states that feature both gas and liquid in a single computational system agree with those from functional line integration for the free energy. Practical implementation of the latter method requires only a cheap numerical routine; the neural functional is the integrand in Eq. (10). The bulk equation of state from functional line integration, Eq. (24), displays a van der Waals loop, see Fig. 6, at subcritical temperatures. We comment further on this result in Sec. IV.3.
@@
The former method of direct numerical stabilization of phase coexistence provides access to one-body liquid-gas interfacial structure. As remarked in Sec. III.4, the density profiles $\rho(x)$ for the two lowest temperatures show the presence of damped oscillations extending into the liquid, see Fig. 7(a). Such behavior is commonly found in DFT calculations that utilize the standard mean-field approximation (4), e.g. Ref. [69] for the square-well and Tschopp et al. [85] for the hard-core Yukawa model. Equivalent behavior is also found in FMT-based DFT calculations for the Asakura-Oosawa model that describes colloid-polymer mixtures. Pronounced oscillations are found on the (colloid-rich) liquid side of the fluid-fluid interface for states in the neighbourhood of the triple point [86]. The physical origin of the oscillations lies, of course, in the packing of particles in the dense liquid. The argument [69], based on asymptotics, is that in the oscillatory region of the bulk phase diagram, as determined by the Fisher-Widom line, the one-body liquid-gas density profile should decay into the bulk liquid at coexistence with the same exponential decay length $1/\alpha_{0}$ and same wavelength $2\pi/\alpha_{1}$ as would be determined for the bulk state from the pole analysis, i.e. Eqs. (17)–(20). Unfortunately there is no simple means of determining the amplitude of oscillations. Moreover, the argument is based on mean-field ideas and omits effects of thermally induced capillary wave fluctuations. The physical picture adopted in Ref. [69], and in subsequent work, is that the DFT results provide a ‘bare’ profile that is then ‘dressed’ by unfreezing the fluctuations which serve to erode the oscillations. For a Gaussian treatment of the fluctuations the decay length and the wavelength of the oscillations are unchanged but the amplitude is reduced by a factor that depends on the interfacial roughness which depends, in turn, upon the interfacial area $A$; see Brader et al. [87] for a thorough discussion. The profiles we find here for the LJ fluid exhibit very weak oscillations even at the lowest temperature shown, suggesting that our neural functional might already capture some fluctuation effects. In this context, it is instructive to take note of Tschopp et al. [85] who show that profiles obtained from a DFT that incorporates correlations into a reference hard-sphere system yields liquid-gas density profiles that have much less pronounced oscillations than those that emerge from employing Eq. (4). It is tempting to surmise that our neural functional provides a very accurate, albeit still mean-field, description of the ‘bare’ interface.
@@
The liquid-gas surface tension is not beset by subtle issues of including capillary wave fluctuations and we expect our results, obtained from an inhomogeneous functional line integral, to be reliable. Indeed these match closely high quality simulation data for the LJ system, see Fig. 7(b). Moreover, our neural functional was able to probe the surface tension for the same range of temperatures as investigated in the calculation of the binodal and we found evidence for a nonclassical critical exponent $\tilde{\mu}$.
@@ IV.3 Discovering phase coexistence and extrapolation of functional mappings
The successful reproduction of results with supercritical training only, cf. Sec. III.7, calls for a reassessment of the importance of the one-body direct correlation functional $c_{1}(\mathbf{r};[\rho],T)$; recall this is the quantity we seek to capture from simulation data via a neural network. It seems feasible to infer this particular functional mapping even when excluding substantial ranges of the parameter space in the training data. Our findings point to the fundamental nature of the one-body direct correlation functional and its favorable mathematical properties. It is arguably the object most appropriate for machine-learning tasks—recall also the simplicity of Eq. (7) for determining $c_{1}(x)$ from simulation data—in contrast to other possible functional mappings [21, 22]. Importantly our results show that liquid-gas phase coexistence is an emerging phenomenon that can already be gleaned far from its onset, i.e. from training data taken above the critical temperature.
@@
We note in this context that the extrapolation capabilities of the neural functional, and of the underlying functional mapping it represents, are arguably much more extensive than one might initially expect. Recall that the data used during training, see Sec. II.3, comprise only true equilibrium states obtained via many-body simulations, and that the resulting density profiles fulfill the Euler-Lagrange equation (8) by construction. However, when using the trained neural functional for predictions, one cannot guarantee a priori that only such ‘physical’ density profiles are encountered. This may be the case during the iteration of Eq. (8) as well as for the evaluation of the excess free energy via the parametrized functional integral (10). For the latter the scaled density profiles $\rho_{a}(\mathbf{r})=a\rho(\mathbf{r})$ are not merely transient, rather they enter as genuine contributions to the total value of $F_{\mathrm{exc}}([\rho],T)$.
@@
That the predictions remain accurate despite requiring the evaluation of the neural functional with possibly ‘unphysical’ density profiles, i.e. ones that cannot correspond to minimization in the presence of any external potential, indicates there is a well-defined continuation of the functional relationship to such states. Furthermore, it seems feasible to infer this extended mapping from physical data alone, using only the restricted function space of the true equilibrium density profiles that provided input. We attribute the successful extrapolation to i) the beneficial mathematical properties of the functional mapping $c_{1}(\mathbf{r};[\rho],T)$ and ii) the prowess of the utilized machine-learning techniques.
This is important in practical applications. Although the evaluation of the neural functional remains valid for ‘unphysical’ density input, the resulting predictions will not correspond directly to what one would find in simulations or in experiments. For instance, the van der Waals loop in Fig. 6 at $T<T_{c}$ arises from the neural prediction for the pressure corresponding to spatially constant bulk densities $\rho(\mathbf{r})=\rho_{b}$, which are imposed to lie in the metastable or unstable coexistence region. Clearly, this is not the sequence of density profiles that one encounters in practice where, for fixed mean density, the system will form liquid and gas domains separated by an interface. If one evaluates the equation of state along this path of physically realizable density profiles, one will indeed find a flat isotherm that is numerically consistent with performing a Maxwell construction in Fig. 6. Of course, this feature is not unique to the neural functional. For an analytical DFT, e.g. employing the mean-field functional (4), one could also solve Eq. (8) for a range of fixed mean densities and determine the resulting density profiles, which will show liquid-gas phase separation within the coexistence region. These will not be as accurate as the present neural results; leave this aside. One could then examine the prediction of the theory for the bulk free energy density and calculate the pressure, analogous to Eq. (24). Making a Maxwell construction will confirm that the coexisting densities agree with those from solving Eq. (8).
The ability to analyze the neural functional for (virtually) arbitrary density input is useful and necessary to make certain calculations. Recall that the determination of the surface tension $\gamma_{lg}$, cf. Fig. 7, requires evaluating the functional line integral (10) along a path of density profiles which crosses the coexistence region. That our neural results for $\gamma_{lg}$ agree well with simulation serves as an indirect verification of the validity of the neural functional predictions for such unseen inputs. Similarly, important bulk quantities such as the isothermal compressibility $\chi_{T}$ and the correlation length $\xi$ can be monitored in states of the bulk fluid that might not be physically accessible, but which provide additional insight—see e.g. the emerging spinodal in Fig. 3. Of course, this is not conceptually different from using analytical DFT: an explicit approximation for the excess free energy functional will provide these quantities for all states. What does differ is the mechanism by which the underlying functional is obtained. Whereas an analytic functional, e.g. Eq. (4), is arguably conceived with a certain phenomenology in mind, features such as a spinodal and a van der Waals loop arise for the neural functional solely from training with appropriately chosen simulation data for a given model fluid.